# The Complexity of Maximizing the MST-ratio

### v1

@@ 2 Maximum MST-ratio of Abstract Graphs
First, we claim that given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value at least $(\ell-2)n+(n-\ell)$ for $G^{\prime}$ in polynomial time. This is trivial if $\ell\leq 2$, and assuming $\ell\geq 3$ it is sufficient to assign the color red to all but one of the vertices in $C$, and blue to the remaining vertices. In this case, the sum of the weights in the MST of the red vertices is $(\ell-2)n$ and for the blue vertices the MST has exactly $(n-\ell)$ many edges and hence the claim follows. Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lceil k/n\rceil}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ be the red MST and blue MST of $G$ in the coloring $Q$. Note that given the weights in $G$, and since $R_{Q}$ and $B_{Q}$ have $n-2$ edges in total, then in total $R_{Q}$ and $B_{Q}$ have exactly $\lceil k/n\rceil$ edges of weight $n$. Without loss of generality, we can assume that $R_{Q}$ has at least $\frac{\lceil k/n\rceil}{2}$ edges of weight. Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight one in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, $r\geq\frac{\lceil k/n\rceil}{2}+1$. For every $1\leq i\leq k$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is the MST on the set of red vertices then for every $i<j$, the weight of the edge $u_{i}u_{j}$ is $n$. Hence $\{u_{1},u_{2},\ldots,u_{k}\}$ is a clique of size $\frac{\ell}{2}+1$ in $G$.
@@
Altogether, this shows that one can use an $\alpha$-approximation algorithm for the MAX-MST-ratio problem to find a $2\alpha$-approximation for the MAX-Clique problem. Moreover, Zuckerman [17] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem unless $P=NP$. Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio. ∎
@@ 3 Maximum MST-ratio of Geometric Point Sets
A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
We rely on a result by Dekster and Wilker [9], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
Lemma 3.
[9, Theorem 2] For every positive integer $n>0$ there exists a real number $10<\lambda_{n}<$ 1 such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
We use the above lemma to increase all the weights of an instance in Theorem 1 by a fixed number in order to make it realizable in Euclidean space.
Lemma 4.
For every weighted complete graph $G$ on $n$ vertices, there exists a number $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
Proof.
Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$. Pick $N$ large enough such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number taken from Lemma 3. Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$. $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$. So $G^{\prime}$ satisfies the conditions of Lemma 3, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$. ∎
Theorem 3.
@@
Proof.
By Lemma 4, the constructed instances in Theorem 1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the minimum spanning trees of the two colors always falls two short of the total number of vertices. Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem. ∎
As established in Theorem 3, determining the maximum EMST-ratio is NP-hard. We conjecture that the problem remains NP-hard even when restricted to a specific dimension. In particular, we are interested in points in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
Conjecture 1.
The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST are NP-hard.
Theorem 2 shows that it is even hard to approximate the maximum MST-ratio for general graphs. However, in Euclidean space, it is not difficult to approximate the maximum EMST-ratio as it cannot be too large. We show a general upper bound for the maximum MST-ratio in any metric space that leads to a simple $3$-approximation algorithm for the MAX-EMST-ratio problem. In the proof, we use the following well-known fact about the path double cover of trees.
Lemma 5 ([15]).
@@
Theorem 4.
In any metric space, the maximum MST-ratio for every set of $n$ points ($n\geq 5$) is not more than $3-\frac{4}{n-1}$.
Proof.
Consider an arbitrary point set $S$ in a metric space, and let $T$ be a minimum spanning tree of $S$. To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ on the set of red vertices, and $B_{Q}$ on the set of blue vertices of $Q$, such that the total weight of $R_{Q}\cup B_{Q}$ is at most $(3-\frac{4}{n-1})|T|$. Let $P^{*}$ be a path with maximum total weight among all paths in $T$ whose endpoints are not leaves of $T$. Assume, without loss of generality, that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$. We traverse the tree from the root $r$ using Depth-First Search (DFS), prioritizing the vertices of $P^{*}$ when exploring branches. The starting time of a node refers to the moment it is first visited during the traversal.
For an arbitrary coloring $Q$ of $S$, we build $R_{Q}$ and $B_{Q}$ in the following way (see Figure 3):
@@
  • Consider any red (blue) vertex $v$ in $T$ and any blue (red) child $w$ of $v$ such that there exists at least one red (blue) vertex in the descendants of $w$ (e.g. $v=r$ in Figure 3). Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$, sorted by starting time (i.e. $u_{1}$ has the earliest starting time, and $u_{x}$ the latest), for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red). We add the edges $u_{1}u_{2},u_{2}u_{3},\ldots,u_{x-1}u_{x},u_{x}v$ to $R_{Q}$ (to $B_{Q}$).
@@
Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 4. The solid black and green edges form the tree $T$ rooted at $r$. The children of any vertex are from left to right according to the order of their starting time in DFS. The green edges are the edges of $P^{*}$.
Let $\mathcal{P}$ be the set of all $uv$-paths $P_{e}$ in $T$, where $e=uv$ is an edge in $R_{Q}\cup B_{Q}$. Observe that every edge of $T$ belongs to at most $3$ paths in $\mathcal{P}$. Namely, by construction, if $x$ is the parent of $y$ and $x$ is blue (similarly red) then the edge $xy$ appears in at most one path $P_{e_{b}}$ where $e_{b}\in B_{Q}$ and at most two paths $P_{e_{r}}$ where $e_{r}\in R_{Q}$. Furthermore, there are certain edges in $T$ that are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e. edges incident to a leaf) and the edges in $P^{*}$ (i.e. the green path in Figure 3).
Since the instance is metric, the weight of every edge $uv$ in $|R_{Q}|\cup|B_{Q}|$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$. So we have $|R_{Q}|+|B_{Q}|\leq 3|T|-W_{L}-W_{P^{*}}$, where $L$ is the set of leaf edges in $T$ and $W_{L}$ and $W_{P^{*}}$ are the total weight of the edges of $L$ and $P^{*}$, respectively. It suffices to show that $W_{P^{*}}+W_{L}\geq(\frac{4}{n-1})|T|$.
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence the claim follows as $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$.
Using Lemma 5, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
@@
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exists a metric space with $n$ points (elements) for which the maximum MST-ratio is $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the points $v_{1},\ldots,v_{n}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to be one. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be zero. We set the weight of any other edge to two. We first argue that the weights are metric. Note that the set of possible weights are $\{0,1,2\}$, the edges of weight one are the edges incident on $v_{n}$, and the edges of weight zero form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight one and has at most one edge of weight zero. Thus $(a,b,c)$ is either $(0,1,1)$, or $(0,2,2)$, or $(1,1,2)$, or $(2,2,2)$, in any case we have $c\leq a+b$. Observe that $|MST(S)|=0\times k+1\times k=k$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red all the other vertices by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$. Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k}=3-\frac{2}{k}=3-\frac{4}{n-1}$.
Theorem 5.
@@
There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof.
@@
By Theorem 4, the maximum EMST-ratio for an $n$-element point set is bounded from above by $3-\frac{4}{n-1}$. On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing the minimum spanning tree of the point set, and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. The running time for computing this coloring is ${O}(n^{2})$, see [8, 16], and clearly, $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$. ∎
We remark that one could achieve a weaker approximation factor $3.47$ in a simpler way, by using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577...}<3.47$ on the MAX-EMST-ratio.
@@
Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 5 can be improved to $2.427+\varepsilon$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [7]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be potential for improvement by discovering more refined colorings in terms of mixedness.
@@
We introduce the Bipartite Coloring as a good candidate to approximate the maximum EMST-ratio of a point set. By bipartite coloring, we mean the partition of a point set into two subsets labeled by colors in which every edge of the minimum spanning tree has different colors on the endpoints. Intuitively, it gives a well-distributed coloring of the point set that can be computed in ${O}(n^{2})$ time. Our experiments suggest that the bipartite coloring gives a very good approximation of the maximum EMST-ratio for random point sets in the plane, see Section 5. Although the bipartite coloring seems to be effective, it does not always produce the maximum EMST-ratio. Here is an extreme example:
@@
Consider $12n+$ 1 points on a zigzag path in the triangular grid and call them a Triangular Chain. Figure 4 illustrates a triangular chain for $n=5$. The maximum EMST-ratio of a triangular chain with $12n+$ 1 points is $\frac{3n-2}{2n}$ while the ratio for the bipartite coloring is $\frac{2n-1}{2n}$. (By stretching the chain slightly, one can ensure a unique MST and, consequently, a unique bipartite coloring.) As $n$ increases, the proportion between the maximum EMST-ratio and the bipartite EMST-ratio approaches $1.5$.
@@
Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of the bipartite coloring. The left panel shows a triangular chain for $n=5$, its EMST in green, and the red and blue EMSTs in the bipartite coloring. The right panel presents the coloring that achieves the maximum EMST-ratio.
@@
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from bipartite coloring for a point set in the plane. Recall that if this conjecture is correct, it would lead to a $O(n^{2})$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 2.
@@
For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for the bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
@@
Apart from the bipartite coloring, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple coloring of a point set.
Question 1.
@@
Is there any polynomial time algorithm for $2$-approximation of the MAX-EMST-ratio problem?
@@
The Average EMST-ratio of a point set $P$ is the average over all $2$-colorings of $P$ of the EMST-ratio. It is automatically a lower bound for the maximum MST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, due to Ondřej Draganov and Morteza Saghafian, the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit. Here, we present the proof for the generalized version of this result, meaning for points within a $d$-dimensional unit cube.
Theorem 6.
@@
For a set of $n$ random points uniformly distributed in $[0,1]^{d}$, the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
Proof.
@@
Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$. Consider all $22^{n}-$ 2 proper bipartitions $P=B\cup R$ into blue and red points. Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bipartitions $|B|=k,|R|=n-k$. Ignoring the relatively small $k$ or $n-k$, by (1), we have
$|EMST(B)|\sim\beta(d).k^{1-1/d},\hskip 14.22636pt|EMST(R)|\sim\beta(d).(n-k)^{1-1/d},\hskip 14.22636pt|EMST(P)|\sim\beta(d).n^{1-1/d}$

### v2

@@ 2 Maximum MST-ratio of Abstract Graphs
Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ denote a red MST and a blue MST of $G$, respectively, in the coloring $Q$. Given the weights in $G^{\prime}$ belong to $\{1,n\}$, and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$. W.l.o.g, we can assume that $R_{Q}$ has at least $\frac{\lfloor k/n\rfloor}{2}$ edges of weight $n$. Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, $r\geq\frac{\lfloor k/n\rfloor}{2}+1$. For every $1\leq i\leq r$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is an MST on the set of red vertices, then for every $i<j$, the weight of the edge $u_{i}u_{j}$ in $G$ is $n$. Hence, $\{u_{1},u_{2},\ldots,u_{r}\}$ is a clique of size $r\geq\frac{\lfloor k/n\rfloor}{2}+1$ in $G$.
@@
Altogether, this shows that one can use an $\alpha$-approximation algorithm for the MAX-MST-ratio problem to find a $2\alpha$-approximation for the MAX-Clique problem. Zuckerman [20] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem, unless $P=NP$. Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio, unless $P=NP$.
@@ 3 Maximum MST-ratio of Geometric Point Sets
We show that MAX-EMST-ratio problem is NP-hard. A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$. We rely on a result by Dekster and Wilker [11], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
Lemma 3.1.
[11, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
We use the above lemma to increase all the weights of an instance in Theorem 2.1 by a fixed number in order to make it realizable in Euclidean space.
Lemma 3.2.
For every weighted complete graph $G$ on $n$ vertices, there exists a positive integer $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
Proof 3.3.
Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$. Pick $N$ such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number in Lemma 3.1. Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$. $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$. So $G^{\prime}$ satisfies the conditions of Lemma 3.1, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$.
Theorem 3.4.
@@
Proof 3.5.
By Lemma 3.2, the instances in Theorem 2.1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the MSTs of the two colors is always two fewer than the total number of vertices. Moreover, the amount we add is $O(n)$, according to Lemmas 3.1,3.2. Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem in polynomial time.
As established in Theorem 3.4, determining the maximum EMST-ratio is NP-hard. We conjecture that the problem remains NP-hard even when restricted to a specific dimension. In particular, we are interested in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
Conjecture 3.6.
The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST-ratio are NP-hard.
Theorem 2.7 shows that it is even hard to approximate the maximum MST-ratio for general graphs. However, in Euclidean space, it is not difficult to approximate the maximum EMST-ratio as it cannot be too large. We show a general upper bound for the maximum MST-ratio in any metric space that leads to a simple $3$-approximation algorithm for the MAX-EMST-ratio problem. In the proof, we use the following well-known fact about the path double cover of trees.
Lemma 3.7 ([18]).
@@
Theorem 3.8.
In any metric space, the maximum MST-ratio for any set of $n\geq 5$ points is at most $3-\frac{4}{n-1}$.
Proof 3.9.
Consider an arbitrary point set $S$ in a metric space, and let $T$ be an MST of $S$. Let $P^{*}$ be a path with maximum total weight among all paths in $T$ whose endpoints are not leaves of $T$. Assume w.l.o.g. that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$. We traverse the tree from the root $r$ using Depth-First Search (DFS), prioritizing the vertices of $P^{*}$ when exploring branches. The starting time of a node refers to the moment it is first visited during the traversal.
To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ and $B_{Q}$ on the set of red vertices and the set of blue vertices of $Q$ respectively, such that $|R_{Q}\cup B_{Q}|$ is at most $(3-\frac{4}{n-1})|T|$. The way we construct $R_{Q}$ and $B_{Q}$ is as follows (see Figure 3).
@@
  • Consider any red (blue) vertex $v$ in $T$ and any blue (red) child $w$ of $v$ such that there exists at least one red (blue) vertex in the descendants of $w$ (e.g. $v=r$ in Figure 3). Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$ for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red). Assume that these vertices are sorted by starting time (i.e. $u_{1}$ has the earliest starting time, and $u_{x}$ has the latest). We add the edges $u_{1}u_{2},u_{2}u_{3},\ldots,u_{x-1}u_{x},u_{x}v$ to $R_{Q}$ (to $B_{Q}$).
@@
Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 3.8. The solid black and green edges form the tree $T$ rooted at $r$. The children of any vertex are from left to right according to the order of their starting time in DFS. The green edges are the edges of $P^{*}$.
Let $\mathcal{P}$ be the set of all $uv$-paths $P_{e}$ in $T$, where $e=uv$ is an edge in $R_{Q}\cup B_{Q}$. Observe that every edge of $T$ belongs to at most $3$ paths in $\mathcal{P}$. Namely, by construction, if $x$ is the parent of $y$ and $x$ is blue (similarly red) then the edge $xy$ appears in at most one path $P_{e_{b}}$ where $e_{b}\in B_{Q}$ and at most two paths $P_{e_{r}}$ where $e_{r}\in R_{Q}$. Furthermore, certain edges in $T$ are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e. edges incident to a leaf) and the edges in $P^{*}$ (i.e. the green path in Figure 3).
Since the instance is metric, the weight of every edge $uv$ in $R_{Q}\cup B_{Q}$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$. So we have $|R_{Q}|+|B_{Q}|\leq 3|T|-W_{L}-W_{P^{*}}$, where $L$ is the set of leaf edges in $T$ and $W_{L}$ and $W_{P^{*}}$ are the total weight of the edges of $L$ and $P^{*}$, respectively. It suffices to show that $W_{P^{*}}+W_{L}\geq(\frac{4}{n-1})|T|$.
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$. Using Lemma 3.7, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
@@
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exist metric spaces with $n$ points (elements) for which the maximum MST-ratio is arbitrarily close to $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the point set $S=\{v_{1},\ldots,v_{n}\}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to be $1$. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be $\varepsilon$ for a sufficiently small $\varepsilon>0$. Any other edge has weight $2$.
We first argue that the weights are metric. The edges of weight $1$ are the edges incident to $v_{n}$, and the edges of weight $\varepsilon$ form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$. Thus $(a,b,c)$ is either $(\varepsilon,1,1)$, or $(\varepsilon,2,2)$, or $(1,1,2)$, or $(2,2,2)$, and in all cases we have $c\leq a+b$. Observe that $|MST(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$. Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
Theorem 3.10.
@@
There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof 3.11.
@@
By Theorem 3.8, the maximum EMST-ratio for an $n$-element point set is upper bounded by $3-\frac{4}{n-1}$. On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [10, 19]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. Finally, as $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
@@
We remark that a slightly improved running time for computing an EMST is provided in [3]. Moreover, one could achieve a weaker approximation factor $3.47$ in a simpler way using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577}<3.47$ on the MAX-EMST-ratio. Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 3.10 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [9]), and for $n>12$ a coloring with EMST-ratio greater than one can be computed in polynomial time (see [13]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be room for improvement by identifying more mingled colorings.
@@
We introduce Bipartite Colorings as good candidates to approximate the maximum EMST-ratio of a point set. By a bipartite coloring of a point set, we mean considering one of the Euclidean minimum spanning trees and then partitioning the point set into two subsets labeled by colors in which every edge of the taken EMST is colorful. Notice that since the EMST of a point set is not necessarily unique, it may have more than one bipartite colorings, namely one for each EMST. Intuitively, bipartite colorings give well-distributed colorings of the point set that can be computed in ${O}(n^{2})$ time ($O(n\log n)$ in the plane, see [5]). Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 5).
@@
Although bipartite colorings seem to be effective, they do not always produce the maximum EMST-ratio. Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain. Figure 4 illustrates a triangular chain for $k=5$. The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$. (By stretching the chain slightly, one can ensure a unique EMST and, consequently, a unique bipartite coloring.) As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
@@
Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of a bipartite coloring. The left panel shows a triangular chain for $k=5$, its EMST in green, and the red and blue EMSTs in a bipartite coloring. The right panel presents the coloring that achieves the maximum EMST-ratio.
@@
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from a bipartite coloring for a point set in the plane. Recall that if this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 3.12.
@@
For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for any bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
@@
Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple mixed coloring of a point set.
Question 3.13.
@@
Does a polynomial-time $2$-approximation algorithm exist for MAX-EMST-ratio?
@@
The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all non-trivial $2$-colorings of $P$. It is automatically a lower bound for the maximum EMST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, it is proved that the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit [12]. We present the proof for the generalized version of this result, i.e. for points within a $d$-dimensional unit cube (or any bounded region in ${{\mathbb{R}}}^{d}$).
Theorem 4.1.
@@
For a set of $n$ random points uniformly distributed in $[0,1]^{d}$ the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
Proof 4.2.
@@
Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$. Consider all $22^{n}-$ 2 proper bi-partitions $P=R\cup B$ into red and blue points. Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bi-partitions $|R|=k,|B|=n-k$. Ignoring the relatively small $k$ or $n-k$, by (1), we know that for every $k_{0}\leq k\leq n-k_{0}$ we have
$|EMST(R)|\sim\beta(d).k^{1-1/d},\hskip 7.96677pt|EMST(B)|\sim\beta(d).(n-k)^{1-1/d},\hskip 7.96677pt|EMST(P)|\sim\beta(d).n^{1-1/d}$